Draw a helical spring of circular wire and -in terms of load, dimensions, and material- express:

  1. - max. shear stress,
  2. - angular deflection,
  3. - spring rate,
  4. - energy stored,
  5. - volume of spring wire,
  6. - axial deflection, and
  7. - length of spring wire.
1
Expert's answer
2022-01-19T19:31:02-0500

helical spring of circular wire


max. shear stress "\tau_{max}= K_s {\frac{8FD}{\pi d^3}}" where "K_s =1+\frac{1}{2C}"

axial deflection, "\delta= \frac{8FD^3N}{Gd^4}"

spring rate, K

length of spring wire. L